Calculate the area of a trapezium that has parallel sides of 10cm and 6cm respectively and the perpendicular distance of 7 cm between the parallel sides.


1
Expert's answer
2021-07-19T05:48:22-0400

"A=\frac{a+b}{2}h=\frac{10+6}{2}*7=56 cm^2."